OpenStudy (hoothoote):

Help! I'll fan and medal! f(x) = 4x + 7, g(x) = 3x2 Find (fg)(x).

OpenStudy (hoothoote):

Answers: 12x2 + 21x 3x2 + 4x + 7 12x3 + 21x2 12x + 21

Nnesha (nnesha):

(fg)(x) is same as f(x) times g(x)

OpenStudy (hoothoote):

would it be A or C?

Nnesha (nnesha):

why do you think its A or C ? please post your work :)

Nnesha (nnesha):

\[f(x) \cdot g(x) \] replace f(x) with (4x+7) and g(x) = 3x^2 and then distribute parentheses by 3x^2

OpenStudy (hoothoote):

Well wouldn't it be 4 times 3? so thats 12x but then I'm confused with the second part and I don't even know if I'm doing it right

OpenStudy (hoothoote):

Ohh wait nevermind

Nnesha (nnesha):

well yeah but don't forget there are variables it is not just 4 and 4 4x times 3x^2

OpenStudy (hoothoote):

Oh okay so is it A?

OpenStudy (hoothoote):

Because the first part would be 12x2

OpenStudy (hoothoote):

?

Nnesha (nnesha):

\[\rm g(x) \cdot f(x) = 3x^2 (4x+2)\] multiply both terms that are inside the parentheses by 3x^2

Nnesha (nnesha):

remember the exponent rule \[\large\rm x^m \cdot x^n =x^{m+n}\] when we multiply same bases we should add their exponents

OpenStudy (hoothoote):

So then the exponents together would be 3?

Nnesha (nnesha):

yes correct !

OpenStudy (hoothoote):

So the the answer would be C??

Nnesha (nnesha):

yes.

OpenStudy (hoothoote):

Thank you!!

Nnesha (nnesha):

np! practice and practice !